OpenStudy (twizttiez):

Jiro dispenses concentrated juice into cylindrical cans from a container that holds 2198 in3. The cans have a radius of 5 in. and a height of 7 in. How many identical cans can Jiro fill? Use 3.14 to approximate pi.

OpenStudy (wolf1728):

Cylinder Volume = PI * r^2 * height Cylinder Volume = 3.14 * 5 * 5 * 7 Cylinder Volume = 549.50 The container contains 2,198 cubic inches which can fill (2,198 / 549.50) or four cans exactly.

OpenStudy (twizttiez):

Okay

OpenStudy (twizttiez):

Ohhhh okay nvm silly me XD

OpenStudy (twizttiez):

I understand

OpenStudy (twizttiez):

Thank you